the area of a trapezoid is 210 square inches. The height is 10 inches and base2 is 24 inches. What is the length of base1?
Answer by checkley71(8403)

210=10(24+x)/2
210=5(24+x)
210=120+5x
5x=210-120
5x=90
x=90/5
x=18 answer for base1.